OpenStudy (anonymous):

What is the y-intercept of the equation 3x - 4y = 16? -4 0 16 3/4

OpenStudy (ivettef365):

Convert the equation in terms of y=mx+b isolate the y, then the result is what you get on b

OpenStudy (ajprincess):

plug in x=0 and find the value of y.

OpenStudy (ajprincess):

that value is the y-intercept:)

OpenStudy (anonymous):

@ajprincess I couldn't do it!

OpenStudy (ajprincess):

3x - 4y = 16 when x=0 3(0)-4y=16 0-4y=16 -4y=16 what will be y?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

right??

OpenStudy (ajprincess):

yup:) good:)

OpenStudy (anonymous):

thansk!

OpenStudy (ajprincess):

welcome:)